Ask
your own question, for FREE!
Mathematics
12 Online
http://prntscr.com/38cj0i
Still Need Help?
Join the QuestionCove community and study together with friends!
Is the correct answer A? -- 2/9?
The probability of having the card B is 1/3 --- {3 cards, 1 chosen} And the probability if rolling a number greater than 2 => -------- case 1 Nos. greater than two = 3,4,5,6 , their probability = 4/6 = 2/3 ---------- case 2 The question asks the prob. of case 1 AND case 2 = p(case1)*p(case 2) = 2.3 * 1/3 = 2/9
I'm guessing the 2.3 is a typo You probably meant to type 2/3 ?
@wolf1728 : Yep :P Sorry, Typo . I meant : 2/3 * 1/3 = 2/9
No problem - just helping out :-)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
5 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal